OpenStudy (anonymous):

Solve the equation using the Zero Product Property. (2x – 4)(2x – 1) = 0

OpenStudy (anonymous):

2x - 4 = 0 2x - 1 = 0 solve each for x

sam (.sam.):

(2x-4)(2x-1)=0 (2x-4)=0 ,(2x-1)=0 ------------------------- (2(x-2))=0 (x-2)=0 x-2=0 x=2 ------------------------- 2x-1=0 2x=1 x=(1)/(2) ====================== x=2,(1)/(2)